### v1

# zkPHIRE: A Programmable Accelerator for ZKPs over HIgh-degRee, Expressive Gates

Abstract
Zero-Knowledge Proofs (ZKPs) have emerged as powerful tools for secure and privacy-preserving computation. ZKPs enable one party to convince another of a statement’s validity without revealing anything else. This capability has profound implications in many domains, including: machine learning, blockchain, image authentication, and electronic voting. Despite their potential, ZKPs have seen limited deployment because of their exceptionally high computational overhead, which manifests primarily during proof generation. To mitigate these overheads, a (growing) body of researchers has proposed hardware accelerators and GPU implementations for kernels and complete protocols. Prior art spans a wide variety of ZKP schemes that vary significantly in computational overhead, proof size, verifier cost, protocol setup, and trust. The latest, and widely used ZKP protocols are intentionally designed to balance these trade-offs. A particular challenge in modern ZKP systems is supporting complex, high-degree gates using the SumCheck protocol. We address this challenge with a novel programmable accelerator that efficiently handles arbitrary custom gates via SumCheck. Our accelerator achieves upwards of $1000\times$ geomean speedup over CPU-based SumChecks across a range of gate types. We integrate this unit into a full-system accelerator, zkPHIRE, which achieves $1486\times$ geomean speedup over CPU and $11.87\times$ speedup over the state-of-the-art at iso-area. zkPHIRE is the first accelerator to scale to problem sizes of $2^{30}$ nominal constraints while maintaining small proof sizes and programmability.
**footnotetext: These authors contributed equally to this work
I Introduction
Zero-Knowledge Proofs (ZKPs) are cryptographic protocols that allow a prover to convince one or more verifiers that a claim is correct without revealing any information beyond the claim’s validity. For example, a prover can demonstrate possession of a valid authentication key without disclosing the key. ZKPs have many compelling applications, including privacy-preserving authentication, trustworthy AI, blockchain scalability, and recently, private age verification [53]. A major limitation of state-of-the-art ZKP protocols is their high computational cost–generating a single proof on a CPU can take minutes to hours. As a result, there is growing interest in accelerating ZKPs, both with programmable solutions (e.g., GPUs [36, 24]) and ASICs [66, 13, 67, 49, 12].
ZKP protocols continue to evolve and improve. Many have been proposed, offering different trade-offs in runtime, proof size, verifier time, and setup (universal vs. application-specific), and security guarantees. A major family of ZKP protocols relies on the SumCheck protocol[35]. SumCheck-based protocols have $O(N)$ prover time, as compared to NTT-based protocols (a popular alternative[20, 17]), which require $O(N\log(N))$ prover operations. The SumCheck protocol is highly flexible and is at the core of many classic and modern ZKP protocols [19, 50, 56, 9, 61, 6, 64], which have relatively fast prover time. While SumChecks were part of previously accelerated systems [55, 49, 12], they studied specific instantiations of the SumCheck protocol rather than the protocol family more generally. Another major distinguishing factor between protocols is the type of polynomial commitment scheme[25] that is used. Some of these protocols, such as Spartan [50]+Orion [61] use faster Merkle tree commitments, at the cost of large proofs (many MBs), while HyperPlonk [9] uses a slower commitment scheme that produces significantly smaller proofs (a few kBs).
ZKP protocols typically encode computations as arithmetic circuits (ACs) – directed acyclic graphs where each node (or gate) enforces a local algebraic constraint, such as addition ($f=a+b$) or multiplication ($f=a\cdot b$) over a finite field. We generally denote the number of gates by $N$. These circuits are then arithmetized by translating gates into polynomial constraints, which are later verified using polynomial checks for correctness. The SumCheck protocol is used to reduce checking $N$ gates to checking a single, randomized gate formula. A key feature of newer SumCheck-based protocols is the use of custom, high degree gates that encapsulate more computation per gate by using higher-degree polynomial constraints (e.g., $f=a\cdot b^{5}+c^{2}$). This reduces the overall number of gates needed to represent a computation, resulting in large efficient gates. The move to custom gates is exemplified in the Halo2 arithmetization language[63]. Halo2 enables a circuit designer to specify the types of gates they want to use in each circuit.
However, this level of customizability introduces significant hardware challenges. Unlike NTTs, which have fixed, regular dataflows, SumCheck computations vary with each gate’s structure and polynomial degree. High-degree polynomials require more modular multiplications and deeper reductions, leading to increased bandwidth pressure, varying memory access patterns, and more complex scheduling. Designing efficient SumCheck hardware thus involves a tradeoff: either build fixed-function units for specific gate shapes, or develop a general, programmable architecture that can dynamically adapt to diverse polynomials without significantly compromising performance.
zkSpeed [12] exemplifies the first approach, using dedicated hardware for Vanilla gates (additions and multiplications) within the HyperPlonk protocol. While this provides good performance for a fixed gate set, it limits flexibility: zkSpeed’s SumCheck unit cannot easily handle complex custom gates like those seen in the Halo2 [63] arithmetization, reducing its applicability across general ZKP workloads. Furthermore, zkSpeed’s reliance on large on-chip scratchpads limits scalability, as its memory footprint grows with the number of gates, becoming impractical beyond $2^{24}$ gates, a scale increasingly common in emerging ZKP applications.
Generalizable SumCheck cores are needed beyond ZKPs, too. In verifiable computing, SumCheck serves as the core verification mechanism in protocols that provide integrity guarantees without requiring zero knowledge. However, previous work in this area, including verifiable ASICs [55], has also relied on custom SumCheck units with fixed functions, again restricting applicability and reuse.
In this work, we adopt the second approach with zkPHIRE, introducing a programmable SumCheck accelerator that supports arbitrary polynomial structures and gate types. zkPHIRE addresses the irregularity and data reuse challenges intrinsic to SumCheck with fused compute pipelines, tree-based interconnects for efficient reductions, and flexible scheduling to handle diverse polynomial structures and degrees. This flexibility opens the door to a wide range of applications across ZKPs and verifiable computing. We evaluate our programmable SumCheck unit on a diverse array of gate types, and then embed it into a full-system accelerator, zkPHIRE. zkPHIRE improves upon zkSpeed with more efficient memory organization and module abstraction that allows hardware to be reused across protocol steps. We conduct a protocol-level evaluation using the HyperPlonk with custom gates as a representative case study. Our key contributions are:
  • A novel, programmable SumCheck unit that supports arbitrary high-degree multilinear polynomials, enabling compatibility with custom gate types.
  • The first hardware evaluation of high-degree gates that offer significant performance and scalability benefits
[…]
  • A full-system accelerator, zkPHIRE, that achieves $11.87\times$ geomean speedup over zkSpeed, $1486\times$ geomean speedup over CPU, and supports succinct proofs for problem sizes up to $2^{30}$ nominal constraints.
II Background
II-A Zero-Knowledge Proofs
Today, the state-of-the-art ZKP protocols are zero-knowledge Succinct Non-interactive Arguments of Knowledge (zkSNARKs). They have three key properties:
  • zero-knowledge: The verifier learns nothing beyond the validity of the statement.
  • succinctness: The proof is short and verified quickly.
  • non-interactivity: The proof is a single message, eliminating the need for back-and-forth communication.
Modern zkSNARKs combine an Interactive Oracle Proof (IOP) and a Polynomial Commitment Scheme (PCS). The IOP encodes the witness and computation as polynomials, which the verifier checks via probabilistic queries. Applying the Fiat-Shamir heuristic [16] makes public-coin IOPs non-interactive. The PCS lets the prover commit to a polynomial and later open it at chosen points, ensuring both binding (the prover cannot change the committed polynomial) and hiding (nothing about the polynomial is revealed until it is opened).
Different zkSNARKs balance proving time, proof size, and security assumptions based on their IOP and PCS designs. Groth16 [20] offers fast verification and tiny proofs, ideal for blockchains, but requires a circuit-specific (i.e., per-application) trusted setup and scales poorly with circuit size. Code-based protocols [61] use error-correcting codes and Merkle commitments to achieve transparency and post-quantum security, but at the cost of multi-megabyte proofs and slower verification. HyperPlonk [9] trades off slightly larger proofs for universal setup and linear-time proving.
II-B Multi-Scalar Multiplication
Multi-Scalar multiplication (MSM) is a key kernel used in polynomial commitments and relies on elliptic curve cryptography. MSMs involve computing a linear combination (i.e., dot product) of scalars and corresponding (2D or 3D) points on an elliptic curve. Given scalars $k_{1},k_{2},\ldots,k_{n}$ and corresponding elliptic curve points $P_{1},P_{2},\ldots,P_{n}$, the goal is to efficiently compute $s=\sum_{i}k_{i}P_{i}$. MSMs are computationally intensive because each term $k_{i}P_{i}$ involves a point-scalar multiplication – a costly elliptic curve operation. MSMs constitute a significant portion of the total runtime in many zkSNARKs, including Groth16 and HyperPlonk. As a result, prior work has focused on accelerating MSMs using both application-specific and programmable platforms [13, 66, 34, 36, 24, 30, 67]. Rather than computing each scalar multiplication directly, optimized approaches such as Pippenger’s algorithm [45] are commonly employed. Pippenger’s algorithm improves efficiency by restructuring the computation to rely more heavily on faster operations like point addition and point doubling.
II-C SumCheck
SumCheck [35] is a round-based Interactive Proof (IP) protocol where a prover, $\mathcal{P}$, attempts to convince a verifier, $\mathcal{V}$, that the sum of a multivariate polynomial, $f(X_{1},\ldots,X_{\mu})$, over the boolean hypercube (all boolean values of $X_{i}\ldots X_{\mu}$) equals a claimed value, $C$. The protocol proceeds in $\mu$ rounds, one for each variable of the polynomial. In each round $i$, $\mathcal{P}$ sends $\mathcal{V}$ a univariate polynomial:
$s_{i}(X_{i})=\sum_{x_{i+1}\in\{0,1\}}\cdots\sum_{X_{\mu}\in\{0,1\}}f(r_{1},\ldots,r_{i-1},X_{i},\ldots,X_{\mu})$
where $r_{1},\ldots,r_{i-1}$ are $\mathcal{V}$’s random challenges from prior rounds. $\mathcal{V}$ checks that the sum of $s_{i}(0)+s_{i}(1)$ equals the value claimed in the previous round (or the initial claim $C$ for the first round). $\mathcal{V}$ then chooses a random field element $r_{i}$ and sends it to $\mathcal{P}$, fixing the variable $X_{i}=r_{i}$ for the next round. After $\mu$ rounds, $\mathcal{V}$ evaluates $f$ at the final point $(r_{1},\ldots,r_{\mu})$. If all checks pass, $\mathcal{V}$ accepts.
ZKPs typically execute the SumCheck protocol over multilinear extensions (MLEs) – polynomials with degree at most 1 in each variable. MLEs can represent multivariate functions compactly and are well-suited for hardware because they can be stored and accessed as flat lookup tables indexed by binary-encoded inputs. Throughout this paper, we will use the terms “MLEs”, “MLE tables,” and “polynomials” interchangeably.
II-C1 Polynomials and Gates
Polynomials are important in ZKP protocols because they encode and help enforce the correctness of the computation being proven. A popular encoding scheme is Plonk [17]. Plonk’s arithmetization scheme converts the underlying computation into an arithmetic circuit consisting of gates. Plonk circuits support additions, multiplications, conditionality, and equality checks. They can be represented by the composite polynomial
$f=q_{L}w_{1}+q_{R}w_{2}+q_{M}w_{1}w_{2}-q_{O}w_{3}+q_{C}$ (1)
consisting of individual polynomials $q_{i}$ and $w_{i}$. These are known as selectors (or enables) and witnesses, respectively. Each of these polynomials encodes a function of a multivariate input vector $X=[X_{1},X_{2},\ldots X_{\mu}]$, which encodes the gate index in binary. For example, if $\mu=3$, there are 8 gates, and gate index 3 is represented by evaluating $f(1,1,0)$. For a multiply operation, we would set $q_{M}(1,1,0)=1$, $q_{O}(1,1,0)=1$ and the other enable polynomials to 0. Then, gate 3 would execute correctly if and only if $f(1,1,0)=0$, that is, $w_{1}(1,1,0)\cdot w_{2}(1,1,0)-w_{3}(1,1,0)=0$. Representing circuits as polynomials naturally suits SumCheck-based provers where we can check (with additional constraints) that each gate evaluates to 0 and that $\sum_{X}f(X)=0$.
II-C2 High-degree, Expressive Gates
The challenge of using Plonk gates is that all computations must be mapped to only what is supported by the Plonk polynomial, similar to mapping program code to a RISC ISA. In some cases, this results in circuits that require millions ($2^{20}$) of gates, and in turn, polynomials with millions of evaluations.
[…]
II-C3 SumCheck on Multilinear Polynomial Products
SumCheck over a single polynomial, such as $\sum f(X_{1},X_{2},X_{3})$ (with $\mu=3$) is relatively straightforward. Using the equation for $s_{1}(X_{1})$ from Section II-C, and viewing $f(X)$ in its MLE table form, in round 1, $s_{1}(0)$ is the sum of all even indices (where $X_{1}=0$), and $s_{1}(1)$ is the sum of all odd indices ($X_{1}=1$). This gives us two evaluations of the univariate, linear polynomial $s_{1}(X_{1})$ to be sent to the verifier. We then perform an MLE Update by fixing $X_{1}$ to random challenge $r_{1}$. At fixed $X_{2}$, $X_{3}$, entries $(0,X_{2},X_{3})$ and $(1,X_{2},X_{3})$ represent a line over $X_{1}$, so fixing $X_{1}=r_{1}$ means extrapolating the line to $r_{1}$: $f(r_{1},X_{2},X_{3})=f(0,X_{2},X_{3})\cdot(1-r_{1})+f(1,X_{2},X_{3})\cdot r_{1}$. This operation halves the size of each MLE table.
SumCheck over polynomials using Plonkish encodings (like Equation 1) is more complicated because there are multiple terms with products of polynomials. Figure 1 shows the dataflow for a single polynomial product term. For a $d$-degree polynomial, we require $d+1$ evaluations for $s_{i}(X_{i})$ after each round of SumCheck. This is done by evaluating all terms at $d+1$ points. Each pair of entries in the MLE table are actually evaluations at $X_{i}=0$ and $1$; these values are then extended to $X_{i}=2,\ldots,d$, using the same formula as MLE Update. Then, the extensions for all polynomials in a term are multiplied together. These products are summed (accumulated) down the MLE table, yielding $d+1$ values for a given term. When there are more than one term, this process will be repeated across terms with the $d+1$ evaluations summed across terms to form the final polynomial $s_{i}(X_{i})$ for round $i$.
III The Programmable SumCheck Accelerator
III-A Challenges
At its core, SumCheck offers many degrees of parallelism. Different MLE tables can compute their extensions in parallel, extensions for different MLE pairs within a table can be computed in parallel, and the products across different extensions can be computed in parallel. Given a SumCheck polynomial, an optimal circuit can be constructed to maximize reuse and throughput, as done in zkSpeed [12].
However, building a generalizable SumCheck unit poses several challenges. For one, the unit must support a diverse set of composite polynomial structures used to express different types of constraints or implement different kinds of operations. The unit must support an arbitrary number of terms and an arbitrary degree. This means that the optimal resource allocation (e.g., number of modular multipliers) for one composite polynomial may be suboptimal or wasteful for another composite polynomial. Additionally, many composite polynomials have individual polynomials that repeat. For example, in $f=a\cdot b\cdot e+c\cdot e+e\cdot g$, the $e$ polynomial appears three times. While we could parallelize all computations for just the first term ($a\cdot b\cdot e$) before proceeding to the next term, this would leave out a clear opportunity for data reuse. Indeed, zkSpeed’s design captures immediate reuse, but is tailored to HyperPlonk’s vanilla gate polynomials. Their design philosophy overprovisions resources for simpler polynomials, and is unable to support complex polynomials.
Figure 1: The SumCheck dataflow of size-$2^{3}$ polynomial $f(X)=abce$. For a degree-4 polynomial $f$, every MLE evaluation pair is used to extend to five evaluations. Extended evaluations are multiplied, summed, and hashed to get a challenge that is used to update MLE tables for the next SumCheck round.
III-B Datapath
Our programmable SumCheck unit seeks to balance the tradeoff between resource allocation and capturing immediate data reuse. Figure 4 shows the datapath for our programmable SumCheck unit. It consists of several processing elements (PEs), with MLE Update units, Extension Engines (EEs) and Product Lanes. Instead of streaming all individual polynomials’ MLE entries simultaneously, we proceed one term at a time, fetching a tile of data from each MLE table and storing in local scratchpad buffers. We allocate 16 scratchpad buffers, more than sufficient to accommodate polynomial structures we see in current ZKP systems. The scratchpad buffers are banked to match the number of PEs for parallel accesses. This ensures that individual polynomials reused in subsequent terms need not be fetched from off-chip memory for a given tile.
For a given term, the polynomials are read into the Update units and EEs, where their extensions are generated by a series of modular adders and subtractors. Specifically, in round 1 of SumCheck, two values are read at a time per MLE and directly fed to generate extensions. In all subsequent rounds, four values are read per MLE, because we pipeline the updates into the extension computation. Updating with four inputs ensures that two values are generated and fed to the EEs to match the throughput of round 1. These updated values are also written to FIFOs that buffer writebacks to off-chip, forming the (halved) tables for the next SumCheck round. In later rounds of SumCheck, when MLE tables can fit fully on-chip, writebacks to off-chip are eliminated and updated values are directly routed to the scratchpad banks.
After the extensions for each MLE are generated, they must be packed into their respective groups, e.g., all extensions at 1 are routed together. Each group of extensions is then fed into a Product Lane (PL). Each PL is equipped with E1E-1 fully pipelined modular multipliers, where $E$ is the number of EEs. Each PL’s output is accumulated into a register corresponding to the extension, e.g., all products at $X_{i}=2$ are always routed to register $2$. We allocate 32 registers (up to degree 31); higher degrees are supported by storing in scratchpads.
III-C Mapping Polynomials to EEs
We devise a scheduler that determines which MLE tables are fed to EEs and PLs using a graph-based decomposition (see Figure 2). In this example, we assume there are 3 EEs, so only 3 MLEs can be handled on-chip at a time. The polynomial is adding a degree-6 term and a degree-3 term. Since there are more MLEs than EEs, we must store the partial products for all extensions; this is done using a dedicated Tmp MLE buffer.
A natural way to schedule the MLEs is using the tree-based approach on the left in Figure 2. However, this approach would require storing two distinct intermediate products that would need to be combined. As the polynomial degree grows, this tree-based approach would require more and more intermediate buffers. Instead, we can use the accumulation based schedule on the right. This scheme only requires 1 Tmp MLE buffer that accumulates extension products within the same term. Notice that this scheme uses the same number of steps as the balanced tree while minimizing temporary storage.
Figure 2: Graph decomposition used to schedule computation on our SumCheck unit. The method on the right minimizes on-chip temporary MLE buffer size. Numbers inside each node indicate the step in the schedule. When processing term hknhkn, $h$ is already prefetched during the prior step.
In this way, the scheduler traverses the polynomial term-by-term, deciding which MLEs are handled by the PEs in each step. These steps are programmed into a controller on-chip. In each step $j$, the PEs read the MLEs identified for step $j$. The controller simultaneously prefetches a tile of data for each MLE in step $j+1$, but only if the MLEs have not already been fetched on-chip. For example, the $h$ MLE would be prefetched during step $2$ (on the right), and the only $k$ and $n$ MLEs would be prefetched during step $3$. Another advantage of this scheme is that prefetch bandwidth is more balanced over steps.
Figure 3: Programmable SumCheck Architecture. The Update unit updates the MLE $a^{\prime}$ to halved-size MLE $a$ for the next SumCheck round. For deg-4 term abceabce, Extension Engines generate five evaluations 0-44.
Figure 4: Overall Architecture of zkPHIRE.
III-D Mapping Extensions to PLs
Since our SumCheck unit must handle any number of extensions, we may run into scenarios where there are $K$ extensions but only $P<K$ lanes. In this case, we have an initiation interval $II=K/P$ and use delay buffers to queue unmatched extensions. In Figure 4, with $K=5$ and $P=3$, the first cycle maps extensions 022 to PLs, while extensions 3344 are buffered for the next cycle. Scheduling then proceeds in a pipelined fashion, with buffered and incoming extensions interleaved across PLs. While $K$ depends on polynomial degree, the schedule for a given $(K,P)$ is static and programmed into the controller at the start of each SumCheck.
III-E Built-in support for ZeroChecks
Many ZKP protocols use SumCheck to verify circuit correctness by checking that a polynomial $f(X)$ evaluates to zero across all inputs. However, simply verifying $\sum_{X}f(X)=0$ is insufficient. Individual gates could evaluate to nonzero values (meaning they were incorrectly executed), but cancel each other out, so the total sum would still be zero. To prevent this, protocols like [50, 61, 7] use ZeroCheck, which multiplies $f(X)$ with a randomized auxiliary polynomial $f_{r}(X)$ and checks that $\sum_{X}(f(X)\cdot f_{r}(X))=0$. $f_{r}(X)$ is constructed on-the-fly from a vector of random challenges generated by a SHA module (referred to as Build MLE in zkSpeed), and zkPHIRE integrates this auxiliary polynomial construction directly into the SumCheck datapath. During the first round, one product lane is dedicated to computing $f_{r}(X)$, with its outputs routed to scratchpad banks and the EEs in parallel with the fetched MLE elements. This reduces the number of lanes for product computations by one in that round but avoids separate $O(N)$ precomputation and memory overhead, reusing existing PL multipliers to mask the cost.
IV zkPHIRE Architecture
We include the programmable SumCheck unit as part of a larger accelerator framework for the HyperPlonk protocol. We describe the steps of the HyperPlonk protocol and its mapping onto zkPHIRE units. HyperPlonk is based on the BLS12-381 elliptic curve, which uses 255-bit scalars (datatypes used by MLEs) and 381-bit elliptic curve points (used for point addition in MSMs). We explore the use of both arbitrary and fixed prime modular multipliers (as done in [49]); the latter saves us roughly 50% on area and improves our computational density by roughly $2\times$. Figure 4 shows the overall architecture.
IV-A HyperPlonk Protocol Overview
HyperPlonk has a very complex dataflow. Here, we describe the high-level protocol and what hardware modules each step requires. The 5 main steps are Witness Commitments, Gate Identity Check, Wire Identity Check, Batch Evaluations, and Polynomial Opening. In Witness Commitments, the prover commits to their private input values (witnesses) using multi-scalar multiplications (MSMs) so that they cannot claim other witness values later. We run this step with the MSM module. In Gate Identity Checks, the prover shows that all gates in the circuit were computed correctly via ZeroChecks. We run this step with our SumCheck unit and a Multifunction Forest unit. In Wire Identity Checks, the prover proves that the wiring between gates is consistent with the circuit’s intended connections. This is done by constructing wiring permutation polynomials and running a PermutationCheck (henceforth called a PermCheck) and then committing to those polynomials. We run this step with a PermCheck Generator unit, the SumCheck and Multifunction Forest units, and the MSM unit. In Batch Evaluations, the prover evaluates all committed polynomials at various points provided by the verifier. We run this step with the Multifunction Forest unit. In the Polynomial Opening, the prover reveals the values of multiple committed polynomials at specific points and proves that it is consistent with the original commitment. The prover then runs a SumCheck to combine these polynomial evaluation claims into one (we call this OpenCheck as per zkSpeed). The combined polynomial commitment is then opened using the MSM unit. Overall this step uses an MLE Combine unit, the SumCheck and Multifunction Forest units, and the MSM unit.
Masking ZeroCheck: In zkSpeed and the HyperPlonk baseline, Gate Identity and Wire Identity steps run serially, despite verifying independent circuit properties. zkPHIRE includes support to overlap the Gate Identity’s ZeroCheck with the MSMs during Wiring Identity. Since MSMs dominate runtime and have low bandwidth pressure due to high data reuse [13, 12], this scheduling effectively masks ZeroCheck latency.
IV-B Accelerator Units
IV-B1 Sparsity Handling for SumCheck
ZKP protocols often exhibit significant sparsity: enable MLEs ($q_{i}$) are mostly binary, while constant and witness MLEs ($q_{c}$, $w_{i}$) are typically 90% sparse [13, 66, 12]. zkPHIRE extends the SumCheck unit to exploit this sparsity for improved efficiency.
zkSpeed opts to use large on-chip scratchpads to store these MLEs, because they are reused throughout steps of the HyperPlonk protocol. However, as the problem size scales, the area consumed by the scratchpad grows, with 10s of MBs dedicated to address translation buffers alone, and a total on-chip storage of nearly 300 MB. Furthermore, the MLEs are only useful for round 1 of SumChecks. For subsequent rounds, the MLEs are updated and written to off-chip memory, so the global scratchpad is idle for roughly 50% of the SumCheck runtime. zkPHIRE opts to use smaller scratchpad buffers to store tiles of each MLE. While this results in more accesses to off-chip memory and we incur more fill/drain latency penalties, it allows for more area to be dedicated core computation structures within SumCheck and other accelerator modules. The improved speedups offset the latency hits incurred by eliminating large scratchpads. To eliminate address translation overheads, zkPHIRE fetches lightweight per-tile offset buffers for witness MLEs. The SumCheck controller decodes 255-bit elements using these offsets, while binary MLEs are stored directly without translation. This scheme incurs minimal bandwidth overhead while enabling more compact storage.
IV-B2 Multifunction Forest
In HyperPlonk, there are several sub-steps that can be mapped to a binary-tree computation, including MLE evaluation, building MLEs, computing the product MLE ($\pi$), and multiplication tree. We adopt the same base tree unit architecture as zkSpeed’s MTU [38]. However, unlike zkSpeed – which employs a single tree unit for all sub-step computations while maintaining a separate, dedicated SumCheck module – zkPHIRE introduces a more unified and efficient approach. In zkPHIRE, we implement the product lanes of the programmable SumCheck unit using a multifunction forest. This allows modular multipliers initially provisioned just for SumCheck to be shared across other protocol steps. The shared multifunction forest optimization helps zkPHIRE achieve the same latency as zkSpeed at iso-workload-length with 15.2% fewer modular multipliers.
IV-B3 MSM
We use the same MSM architecture as zkSpeed, since we target the same protocol, HyperPlonk. In HyperPlonk, the number of MSMs only depends on the number of witnesses. The underlying hardware is unchanged, we just compute more MSMs using it. Jellyfish gates have five witnesses, so we run five Sparse MSMs and three Dense MSMs.
IV-B4 MLE Combine
The MLE Combine module, shown in Figure 4, is used for various element-wise operations and dot products before and after the OpenCheck in Polynomial Opening. For these steps, we fetch MLE tables from off-chip into up to 6 local SRAM buffers and perform the operations between MLE entries and/or challenges from prior protocol steps stored in registers. All operations are fully pipelined.
IV-B5 PermCheck Generator
Figure 5: The PermCheck Generator Module.
PermCheck requires constructing Numerator (N), Denominator (D), Fraction ($\phi$), and Product ($\pi$) MLEs [12]. Since $\phi$ is computed on-the-fly from N and D, we design a pipelined unit that generates all three in parallel, producing one element per cycle after warmup (Figure 5). The Multifunction Forest is used to construct $\pi$ from $\phi$, as in zkSpeed. To accommodate both vanilla and custom gates, we allocate 5 PEs (one per witness), with the ability to support more than 5 through overlapped scheduling and cyclic reuse of PEs. Intermediate $N_{1\dots k}$ and $D_{1\dots k}$ MLEs are written to HBM and combined via modular multiplications to produce the final N and D MLEs, followed by modular inversion of D.
[…]
IV-B6 Interconnect and Memory
HyperPlonk’s stage-level data obliviousness enables static scheduling of computation and communication via a centralized controller. zkPHIRE comprises six major modules connected by a multi-channel shared bus, which is provisioned to meet peak data movement demands. During SumCheck and Wiring Identity, where bidirectional transfers occur between the SumCheck unit and Multifunction Forest, and one-way from PermCheck Generator to MSM unit, up to three channels are needed to avoid stalls. At our 294mm2 design point, the peak bandwidth requirement reaches 19 TB/s, comparable to state-of-the-art accelerators  [47, 28].
Each module includes local SRAM scratchpads to buffer intermediate data. Smaller buffers (6 MB) serve PermCheck Generator, MLE Combine, and the Multifunction Forest. Heavier modules—MSM (43 MB) and SumCheck (6 MB)—use highly banked SRAM to support multiple PEs and exploit data reuse. For instance, MSM stores elliptic curve points, while SumCheck stores frequently accessed MLE tiles. All modules connect to off-chip PHYs via the shared bus, which is coordinated to avoid contention and sustain peak throughput.
V Methodology
[…]
The HyperPlonk CPU performance is measured on an AMD EPYC 7502 32-core processor running up to 3.35GHz, and the total die size is 296 mm2 [41, 5, 54]. We benchmark our SumCheck unit performance relative to prior work [12] and a 4-threaded CPU implementation. We then use workloads from prior work [9, 1] to model the performance of our full architecture vs. 32 CPU threads, up to problem sizes of $2^{30}$ nominal gates. We show these results in Section VI. We use a cycle-accurate simulator to model runtimes for MSMs. For SumCheck and tree-based computations, we model performance using analytical models, since these steps are data-oblivious. For sparse computations, we use the same workload statistics from prior work [13, 36, 66, 1].
VI Evaluation
We present our evaluation in two parts. First, we evaluate our programmable SumCheck unit as a standalone unit in Section VI-A. We use polynomial gates seen in prior works [55, 50] as well as gates used for elliptic curve addition in the Halo2 [63] library. We then evaluate it as part of a broader HyperPlonk accelerator, zkPHIRE, in Section VI-B.
Table I: List of polynomial constraints.
ID Name Polynomial
0 Verifiable ASICs [55] $q_{\text{add}}\cdot(a+b)+q_{\text{mul}}\cdot(a\cdot b)$
1 Spartan 1 [50] $(A\cdot B-C)\cdot f_{\tau}$
2 Spartan 2 [50] $(\text{Sum}_{ABC})\cdot Z$
Halo2 Constraints [63]
3 Nonzero Point Check $q_{\text{point}}^{\text{non-id}}\cdot(y^{2}-x^{3}-5)$
4 $x$-gated Curve Check $(q_{\text{point}}\cdot x)\cdot(y^{2}-x^{3}-5)$
5 $y$-gated Curve Check $(q_{\text{point}}\cdot y)\cdot(y^{2}-x^{3}-5)$
6 Incomplete Addition 1
$q_{\text{add-incomplete}}\cdot((x_{r}+x_{q}+x_{p})$
$\cdot(x_{p}-x_{q})^{2}-(y_{p}-y_{q})^{2})$
7 Incomplete Addition 2
$q_{\text{add-incomplete}}\cdot(y_{r}+y_{q})(x_{p}-x_{q})$
$(y_{p}-y_{q})(x_{q}-x_{r})$
8 Complete Addition 1
$q_{\text{add}}\cdot(x_{q}-x_{p})((x_{q}-x_{p})\lambda$
$(y_{q}-y_{p}))$
9 Complete Addition 2
$q_{\text{add}}\cdot(1-(x_{q}-x_{p})\alpha)$
$(2y_{p}\lambda-3x_{p}^{2})$
10 Complete Addition 3
$q_{\text{add}}\cdot x_{p}x_{q}(x_{q}-x_{p})$
$(\lambda^{2}-x_{p}-x_{q}-x_{r})$
11 Complete Addition 4
$q_{\text{add}}\cdot x_{p}x_{q}(x_{q}-x_{p})$
$(\lambda(x_{p}-x_{r})-y_{p}-y_{r})$
12 Complete Addition 5
$q_{\text{add}}\cdot x_{p}x_{q}(y_{q}+y_{p})$
$(\lambda^{2}-x_{p}-x_{q}-x_{r})$
13 Complete Addition 6
$q_{\text{add}}\cdot x_{p}x_{q}(y_{q}+y_{p})$
$(\lambda(x_{p}-x_{r})-y_{p}-y_{r})$
14 Complete Addition 7 $q_{\text{add}}\cdot(1-x_{p}\cdot\beta)(x_{r}-x_{q})$
15 Complete Addition 8 $q_{\text{add}}\cdot(1-x_{p}\cdot\beta)(y_{r}-y_{q})$
16 Complete Addition 9 $q_{\text{add}}\cdot(1-x_{q}\cdot\gamma)(x_{r}-x_{p})$
17 Complete Addition 10 $q_{\text{add}}\cdot(1-x_{q}\cdot\gamma)(y_{r}-y_{p})$
18 Complete Addition 11
$q_{\text{add}}\cdot(1-(x_{q}-x_{p})\alpha$
$(y_{q}+y_{p})\delta)x_{r}$
19 Complete Addition 12
$q_{\text{add}}\cdot(1-(x_{q}-x_{p})\alpha$
$(y_{q}+y_{p})\delta)y_{r}$
HyperPlonk Polynomials [9]
20 Vanilla ZeroCheck
$(q_{L}w_{1}+q_{R}w_{2}-q_{O}w_{3}$
$q_{M}w_{1}w_{2}+q_{C})f_{r}$
21
Vanilla PermCheck
($\alpha$ is a scalar)
$(\pi-p_{1}p_{2}+\alpha(\phi D_{1}D_{2}D_{3}$
$N_{1}N_{2}N_{3}))f_{r}$
22 Jellyfish ZeroCheck
$(q_{1}w_{1}+q_{2}w_{2}+q_{3}w_{3}+q_{4}w_{4}$
$q_{M_{1}}w_{1}w_{2}+q_{M_{2}}w_{3}w_{4}$
$q_{H_{1}}w_{1}^{5}+q_{H_{2}}w_{2}^{5}+q_{H_{3}}w_{3}^{5}$
$q_{H_{4}}w_{4}^{5}-q_{O}w_{5}$
$q_{ecc}w_{1}w_{2}w_{3}w_{4}+q_{C})f_{r}$
23
Jellyfish PermCheck
($\alpha$ is a scalar)
$(\pi-p_{1}p_{2}+\alpha(\phi D_{1}D_{2}D_{3}\cdot$
$D_{4}D_{5}-N_{1}N_{2}N_{3}N_{4}N_{5}))f_{r}$
20 OpenCheck
$y_{1}f_{r_{1}}+y_{2}f_{r_{2}}+y_{3}f_{r_{3}}$
$y_{4}f_{r_{4}}+y_{5}f_{r_{5}}+y_{5}f_{r_{5}}$
Figure 6: Speedups over 4-threaded CPU implementations of SumChecks on the polynomials in Table I. We pick a different optimal design across each bandwidth technology. Utilization across the design points shown
VI-A Programmable SumCheck
VI-A1 Speedup over various gate types
Table I enumerates the polynomials we use in our evaluation. We first benchmark the performance of SumChecks on each polynomial by extending HyperPlonk’s existing code base [9]. We then run each polynomial on our programmable SumCheck unit. In this evaluation, we seek to optimize both performance and utilization across all polynomials. We can always allocate EEs and PLs to handle the worst-case polynomial degree; however, this would lead to underutilization on low-degree polynomials. In our design space search, we specify an area and bandwidth constraint and then use the following objective function:
$\min_{\text{design}}\;\lambda\cdot(1-f_{\text{util}}(u_{d,i}))+\lambda\cdot f_{\text{slowdown}}(s_{d,i})$
Here, $s_{d,i}$ refers to the slowdown of the $i^{th}$ polynomial for a given design $d$, relative to the fastest runtime for that polynomial in the area-constrained design space. Similarly, $u_{d,i}$ reflects the utilization for design $d$ on polynomial $i$. $f_{\text{slowdown}}$ and $f_{\text{util}}$ are aggregation functions; we employ geomean for slowdown and arithmetic mean for utilization. $\lambda$ controls the tradeoff between the minimizing slowdown and maximizing utilization. For this analysis, we focus on utilization, using $\lambda=0.8$. We compare the runtime against the CPU running each polynomial’s SumCheck on 4 threads, for which we estimate the total core area to be 37 mm2 in 7nm [41, 5, 54] and use that as our area constraint (after scaling from 22nm). Across various bandwidths, we find the optimal design point and report speedups and utilizations in Figure 6. For most designs, two EEs and five PLs is optimal.
We observe  50% utilization across benchmarks at each bandwidth, due to three main factors: (1) MLE Update units are inactive during the first round of SumCheck, which accounts for about half of total runtime; (2) lower-degree polynomials require fewer product lanes; and (3) repeating MLEs (e.g., $q_{add}$) reduce update modmul demand since they are reused across terms. Consequently, when optimizing for utilization, we observed a tendency to limit the number of EEs. Further parallelism and potentially higher utilization can be achieved by mapping multiple terms to EEs when possible, but this would introduce additional complexity in interconnect and increase bandwidth demand. These tradeoffs can be explored further in future work. In contrast, product lane modmuls are usually fully utilized, provided the polynomial degree and memory bandwidth are high enough. Despite moderate utilization, we achieve significant speedups over CPU baselines, reaching nearly 1000× at 1 TB/s of bandwidth.
VI-A2 High degree sweep
Figure 7: Performance of a fixed SumCheck configuration on high-degree polynomials at different bandwidths. Speedup relative to CPU is also plotted.
We analyze the performance scaling of SumChecks as we increase polynomial degree. We sweep polynomial degrees ($d=2\ldots 30$) for $f=q_{1}w_{1}+q_{2}w_{2}+q_{3}w_{1}^{d-1}w_{2}+q_{c}$, as in HyperPlonk, and measure speedup over CPU. We use the same objective as before and pick a high-performance design under same area constraints (Figure 7). Our results show that low-degree polynomials require HBM-scale bandwidths to achieve $1000\times$ speedups, while higher-degree polynomials can reach similar speedups with DDR5-level bandwidths (e.g., 256 GB/s). Higher-degree polynomials in this scenario exhibit more computations on the same amount of data, which is why the disparity in speedups across bandwidths is less compared to low-degree polynomials.
Figure 8: Scheduler-induced behavior of SumCheck at fixed bandwidth and PL settings.
At fixed bandwidth, PE count, and EE-per-PE settings, increasing the number of PLs per PE improves performance roughly proportionally. Similar gains are observed when increasing EE count at fixed PLs. Interestingly, as shown in Figure 8, runtime increases in discrete jumps due to the graph decomposition (see Figure 2) the scheduler uses to map polynomials to EEs. For example, under 6 EEs, degree-1166 polynomials have 1 node in their schedule graph, while degree-771111 polynomials require 2. Each added node incurs a sharp runtime jump, as it introduces additional products across all MLEs and terms. Within each node cluster, runtime grows more gradually with degree due to early exit optimizations. These trends may help optimize the SumCheck unit when the degree distribution of polynomials is known, by minimizing high-degree outliers that cause runtime spikes.
VI-A3 Comparison with Prior ASICs
Figure 9 details our speedups relative to a prior ASIC, zkSpeed. zkSpeed implements the SumChecks by using a custom unified core that maximizes immediate reuse. We also compare with zkSpeed+, which gives zkSpeed the benefits of pipelining MLE updates directly into the extensions and product computations (as we do in zkPHIRE). We assume the same modular multiplier area and 2 TB/s as zkSpeed. We use the same objective function from Section VI-A1 on the same “training set” of polynomials in Table I with a roughly iso-zkSpeed area constraint. We pick a 35.24 mm2 design to compare with zkSpeed’s 30.8 mm2 SumCheck and MLE Update area. Our design has 4.870 mm2 of SRAM, while zkSpeed used a much larger global SRAM for storing MLEs, hence we believe this comparison is fair.
We run the same polynomials from zkSpeed’s evaluation on our SumCheck unit, and find that compared to the zkSpeed+, we are only 30% slower at iso-area and iso-bandwidth, while offering programmability to support other types of polynomials beyond those in HyperPlonk. We then use the same hardware design point for evaluation on Jellyfish polynomials. The advantage of using Jellyfish gates is that for a given application, while the polynomial complexity grows, the total gate count (i.e., workload length) decreases. We simulate three degrees of gate count reduction, $2\times,4\times,\text{and }8\times$.
We find that, for ZeroCheck, because the Jellyfish polynomial is more complex, the $2\times$ is not sufficient to bring speedups over the vanilla counterpart. However, PermCheck’s increase in complexity is outweighed by a $2\times$ workload reduction. OpenCheck, which is the same between Jellyfish and Vanilla, exhibits $1.12\times,3.54\times$ and $6.22\times$ speedup for each gate reduction over zkSpeed+. Combining all three SumChecks, a Jellyfish gate count reduction by $4\times$ is sufficient to outperform vanilla polynomials on both zkSpeed+ and zkPHIRE. These gains are observed despite the inherent overheads of programmability in our accelerator.
Figure 9: Comparison with prior ASICs. Bars are annotated with speedup numbers over baseline zkSpeed / zkSpeed+.
VI-B zkPHIRE
We now incorporate the programmable SumCheck unit into a larger accelerator that accelerates the HyperPlonk protocol, zkPHIRE. Here, our analysis focuses on yielding high-performance designs. In these comparisons, we assume the CPU runs with 32 threads, roughly 296 mm2 in 7nm. These experiments assume fixed-prime modular multipliers, which is a common technique seen in prior cryptographic accelerators and libraries [49, 9, 1].
Table II: Design Space Exploration Parameters for zkPHIRE.
Module Design Knob Values
SumCheck PEs 1, 2, 4, 8, 16, 32
SumCheck Extension Engines 2, 3, 4, 5, 6, 7
SumCheck Product Lanes 3, 4, 5, 6, 7, 8
SumCheck SRAM Bank Size $2^{10}-2^{15}$
MSM PEs 1, 2, 4, 8, 16, 32
MSM Window Size 7, 8, 9, 10
MSM Points/PE 1K, 2K, 4K, 8K, 16K
FracMLE PEs 1, 2, 3, 4
Bandwidth (GB/s) 64, 128, 256, 512, 1T, 2T, 4T
Figure 10: Pareto Frontiers for $2^{24}$ gates. We plot the individual Pareto curves for each bandwidth tier and the global Pareto frontier in the inset.
VI-B1 Pareto Frontier Analysis
We sweep the parameters in Table II to obtain Pareto-optimal designs for each bandwidth, and then construct a global Pareto frontier across all bandwidths. Figure 10 illustrates the design space for a $2^{24}$ Jellyfish gate workload under seven bandwidth scenarios, ranging from DDR-class to HBM-scale bandwidths. We account for memory PHY overheads [49, 15, 39, 22, 47, 48, 27, 26], assuming 14.9 mm2 per HBM2 PHY and 29.6 mm2 per HBM3 PHY [2]. For $2^{24}$ Jellyfish gates, the CPU runtime is roughly 182.896 seconds. We can hit roughly $1000\times$ speedup with a 207 mm2 design and 1 TB/s off-chip bandwidth, and with $294$ mm2 and 2 TB/s, around $1400\times$ at iso-CPU area. For this high-performance regime, a Pareto design around 250 mm2 with 1 TB/s (design B) achieves nearly $2\times$ and $3\times$ speedup over 512 GB/s and 256 GB/s configurations. These results highlight a key trend in recent cryptographic accelerators where high performance demands investments in high-bandwidth memory technologies (even for small accelerators like NoCap, where the PHY cost alone is $3\times$ the compute area). That being said, with our accelerator framework, for applications with looser prover constraints, sub-DDR-scale bandwidths can still yield $100-200\times$ speedups over multi-threaded CPU baselines.
Figure 11: Area (left, top legend) and runtime (right, bottom legend) breakdowns for the selected Pareto points in Figure 10.
Figure 12: Runtime breakdown for CPU and zkPHIRE for $2^{24}$ Jellyfish gates. CPU’s sequential execution enables finer breakdowns compared to zkPHIRE.
VI-B2 Area, Runtime, and Power
In Figure 11, we analyze selected Pareto points from Figure 10. Across all points, MSM dominates the area as expected. As the bandwidth increases, both MultiFunction Forest and MSM portions increase, but from C to D, the absolute MSM area remains unchanged while the SumCheck and Forest area increase. This is also evident in the runtime breakdown: from C to D, the SumCheck portions (ZeroCheck, PermCheck, and OpenCheck) become smaller. Because SumCheck is memory-bound, higher bandwidth incentivizes more resource allocation towards SumCheck-related compute. For low-performance designs, less bandwidth is required, so the MSM is allocated a higher area portion.
Table III: Area and power of zkPHIRE. Other includes the PermCheck Generator, MLE Combine and SHA3 units.
Area (mm²) Average Power (W)
MSM (32 PEs) 105.69 58.99
Multifunc Forest (80 Trees) 48.18 40.69
SumCheck (16 PEs) 16.65 14.43
Other 10.64 6.17
Total Compute 181.15 120.29
SRAM 27.55 3.56
Interconnect 26.42 14.83
HBM3 (2 PHYs) 59.20 63.60
Total Memory System 113.18 81.99
Total 294.32 202.28
Table III shows an exemplar zkPHIRE design point we choose for the following analysis. In this 294 mm2 design, there are 32 MSM PEs, with 8 modular multipliers per Multifunction tree PE, and 7 EEs and 5 PLs per SumCheck PE. The design runs on 2 TB/s HBM bandwidth. We use two $\times32$ bit-sliced crossbars [47, 28, 44] in the MSM and SumCheck unit, with a shared-bus to form the on-chip bandwidth (up to 19 TB/s) between modules. Figure 12 further breaks down the runtime to compare with HyperPlonk on CPU. The colors indicate the same HyperPlonk step between CPU and zkPHIRE running at the configuration of Table III. We show the latency breakdown for Gate and Wire Identities before performing ZeroCheck masking to show their initial runtime proportions. MSMs dominate runtime before and after acceleration, though compared to zkSpeed’s CPU baseline [12], SumChecks take more runtime because the polynomials are complex, exerting more memory pressure on the CPU.
Figure 13: zkPHIRE speedups across workloads relative to Vanilla Gates. MskZC = Masked ZeroCheck optimization.
Figure 14: High-degree sweep on overall HyperPlonk Protocol. Crossover point is noted. $f=q_{1}w_{1}+q_{2}w_{2}+q_{3}w_{1}^{d-1}w_{2}+qc$.
VI-B3 Performance from Protocol-level Optimizations
Figure 13 quantifies the benefits of using Jellyfish gates and masking ZeroCheck latency across several workloads, since different workloads may respond differently to Jellyfish-style mappings. For large workloads, where the Jellyfish gate count is $2^{20}$, the speedups from Jellyfish alone reasonably reach what we expect from table size reduction alone. For smaller workloads, we see lower speedups from Jellyfish alone due to serialization overheads of MSMs and other off-chip memory overheads (e.g. fill/drain latencies) that are dwarfed for large workloads. To highlight this, we test scaled up versions of Zcash and Zexe to $2^{24}$ and $2^{25}$ (as done previously [49]), and observe $3.1\times$ and $23.3\times$ speedups from Jellyfish alone. For zkEVM, no estimation of vanilla gates is available [9]; we estimate an $8\times$ reduction to highlight our speedup trends for a hypothetical case. ZeroCheck masking provides additional gains of roughly 25-27% for most workloads, which is what we expect from Amdahl’s law by inspecting Figure 12.
VI-B4 High-Degree Sweep
In Figure 14, we plot the runtime of the highlighted design point over custom gates represented by $f=q_{1}w_{1}+q_{2}w_{2}+q_{3}w_{1}^{d-1}w_{2}+qc$. For this family of gates, the number of witnesses is fixed, so the total MSM runtime is constant across all values of $d$. Consequently, there is a crossover point after which SumChecks dominate the total protocol runtime. This highlights a key design tradeoff: while higher-degree gates reduce overall gate count, they can potentially shift the performance bottleneck from MSMs to SumChecks. Protocol designers must therefore balance gate expressivity against the rising cost of polynomial verification.
VI-B5 Performance Comparison
Table IV compares a 300 mm2 zkPHIRE design using the same modmuls as zkSpeed+ (366 mm2 with MLE Updates pipelined into SumCheck) and without ZeroCheck masking. This is to provide as fair a comparison as possible (zkSpeed+ is roughly $10\%$ faster than zkSpeed). Speedups are reported relative to CPU. zkPHIRE is about $10\%$ slower than zkSpeed+, while offering flexibility to support arbitrary, custom gates. Table V shows zkPHIRE’s speedups at iso-CPU area (294 mm2) with fixed primes and ZeroCheck masking. We achieve $1486\times$ geomean speedup and, for the first time, bring workloads on the order of $2^{30}$ nominal gates within reach for proof generation. Lastly, in Table VI, we compare the same design (with ZeroCheck masking and fixed primes) with zkSpeed+ at iso-application, achieving geomean $11.87\times$ speedup.
Table IV: Comparison with zkSpeed+ and CPU for Vanilla gates.
Workload Gates Runtime (ms)
CPU zkSpeed+ zkPHIRE
[…]
Table V: Comparison with CPU for Jellyfish gate runtime. Vanilla gate counts are shown to show original problem size.
Workload Gates Runtime (ms)
Vanilla Jellyfish CPU zkPHIRE
[…]
Table VI: Iso-application comparison of zkSpeed+ (Vanilla) and zkPHIRE (Jellyfish).
Workload Gates Runtime (ms)
Vanilla Jellyfish zkSpeed+ zkPHIRE
[…]
VII Related Work
Table VII: Comparison of zkPHIRE with Prior ZKP Accelerators. N = NTT, S = SumCheck, M = MSM. Areas are in 7nm.
Metric NoCap SZKP+ zkSpeed+ zkPHIRE
Workload Scaled-up AES
Rollup 25
Pvt Tx
Rollup 25
Pvt Tx
Rollup 25
Pvt Tx
Protocol Spartan+Orion Groth16 HyperPlonk HyperPlonk
Main Kernels N & S N & M S & M S & M
Gates $^{24} 22$$^{24}$ $2^{24}$ $2^{19}$
Encoding R1CS R1CS
Plonk
Plonk
Proof Size 8.1 MB 0.18 KB 5.09 KB 4.41 KB
Setup none circuit-specific universal universal
Prime fixed arbitrary arbitrary fixed
Bitwidth 64 255/381 255/381 255/381
SW Prover (s) 94.2 51.18 145.5 6.161
HW Prover (ms) 151.3 28.43 151.973 3.874
SW Verifier (ms) 134 4.2 26 19
Chip Area (mm2) 38.73 353.2 366.46 294.32
# Modmuls 2432 1720 1206 2267
Modmul (mm2) 0.002 0.133 / 0.314 0.133 / 0.314 0.073/0.162
Power (W) 62 >>220 171 202
[…]
Nocap [49] implements SumChecks using their vector architecture and chip. While SumCheck is highly parallelizable, its core operation—reducing a vector—can require repeated data access and movement on a vector processor. Reductions over a vector of length $M$ require $\log_{2}(M)$ folding steps, typically implemented as shift-and-add instructions, that must be serialized and access the vector register file. Furthermore, higher degree polynomials introduce more extensions, each of which requires a dedicated vector reduction, causing the total reduction overhead grows with degree. In contrast, zkPHIRE uses fused pipelines with tree-like, direct connections to support efficient reductions with minimal data movement. Modern ZKP protocols rely on distinct kernels like SumCheck and MSM, which differ in compute and communication patterns. zkPHIRE’s heterogeneous architecture provides specialized hardware for each, avoiding the inefficiencies of a general vector-based design balanced to handle both.
VIII Conclusion
This paper presents zkPHIRE, the first accelerator to generate succinct proofs for problem sizes reaching $2^{30}$ nominal constraints. zkPHIRE supports general, programmable SumCheck, enabling ZKPs on custom, high-degree gates. zkPHIRE outperforms state-of-the-art accelerators for HyperPlonk, achieving iso-application geomean speedups of $11.87\times$ over ASIC and $1486\times$ over CPU.